OpenStudy (anonymous):

I need help with problem set 5, 2C-1d. I think my answer is right but I don't know how to simplify it so it looks like the solution. Then again, my answer could be completely wrong.

OpenStudy (anonymous):

find dw for \[w = \sin^{-1}( u/t)\] bearing in mind that if \[f(x) = \sin^{-1} (x)\] then \[f'(x) = 1/\sqrt{1 - x^2}\] and the product rule, then my solution is \[1/\sqrt{1-(u/t)^2}.du/t + 1/\sqrt{1-(u/t)^2}.u.dt\] which simplifies to \[ (du/t + u.dt)/\sqrt{1-(u/t)^2}\] The answer provided is \[(tdu - udt)/t.\sqrt{t^2-u^2}\] Where am I going wrong?

OpenStudy (anonymous):

\[w = \sin ^{-1}(u/t)\] \[\sin (w)=(u/t)\] \[\cos (w)dw=\frac{1}{t}du-\frac{u}{t^2}dt\] since \[\cos (w)=\sqrt{1-\sin ^2(w)}\] and \[\sin ^2(w)=\frac{u^2}{t^2}\] then \[\cos (w)=\sqrt{1-\frac{u^2}{t^2}}\] now, substitute that into the above equation and you get \[\ \sqrt{1-\frac{u^2}{t^2}} dw=\frac{1}{t}du-\frac{u}{t^2}dt\] \[dw=\frac{(\frac{1}{t}du-\frac{u}{t^2}dt)}{\sqrt{1-\frac{u^2}{t^2}}}\] \[dw=\frac{(\frac{t}{t^2}du-\frac{u}{t^2}dt)}{\sqrt{1-\frac{u^2}{t^2}}}\] \[dw=\frac{(tdu-udt)}{t^2\sqrt{1-\frac{u^2}{t^2}}}\] \[dw=\frac{(tdu-udt)}{t\sqrt{t^2-u^2}}\]